Project

General

Profile

NDOS Commissioning Meeting 2012/01/19

Attendance: Jon, Athans, Jan, Minerba, Leon, Mat, Denis, Andrew, Martin

Agenda:

  1. Beam and Detector Update
  2. Installation Plans for the next week
  3. DAQ Update and Plans for next week
  4. DCS Update and Plans for next week
  5. AOB

Beam and Detector Report

  • Jon's Report: doc-dB 6986
  • In general the detector showed no significant issues this past week. The beam was off for most of yesterday. DAQ commissioning studies were done during much of Tuesday and Wednesday.

Installation plans for next week

  • Mat: may be doing some tests with the APD gas-drying system.

DAQ Update and Plans for next week

  • No report.
  • Jon: the novadbprod dB machine is scheduled to be down for most of next Thursday for an OS upgrade. Jon is working on migrating the dB tables from that server to a DAQ-owned dB server. Hopefully this will be worked out by early next week. Replication of the DAQ-owned dB tables to the offline dB will likely be done by brute force until a more elegant solution is found.

DCS Update and Plans for next week

  • Nothing specific to report.

AOB

  • None.